OpenStudy (anonymous):

Evaluate 1\frac 14\cdot 4\frac 45. Express your answer in simplest form.

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{14}\cdot \cfrac{4}{45}\implies \cfrac{1}{7\cdot \cancel{ 2 }}\cdot \cfrac{2\cdot \cancel{ 2 }}{45}\)